# High Redshift Merger Model for Low Frequency Gravitational Wave Background

### v1

@@ Abstract
In 2023, the Pulsar Timing Array (PTA) Collaborations announced the discovery of a gravitational wave background (GWB), predominantly attributed to supermassive black hole binary (SMBHB) mergers. However, the detected GWB is several times stronger than the default value expected from galactic observations at low and moderate redshifts. Recent findings by the James Webb Space Telescope (JWST) have unveiled a substantial number of massive, high-redshift galaxies, suggesting more massive SMBHB mergers at these early epochs. Motivated by these findings, we propose an “early merger” model that complements the standard merger statistics by incorporating these early, massive galaxies. We compare the early and standard “late merger” models, which assume peak merger rates in the local universe, and match both merger models to the current detected GWB. Our analysis shows that the early merger model has a significantly lower detection probability for single binaries and predicts a $\sim 30\%$ likelihood that the first detectable single source will be highly redshifted and remarkably massive with rapid frequency evolution. In contrast, the late merger model predicts a nearly monochromatic first source at low redshift. The future confirmation of an enhanced population of massive high-redshift galaxies and the detection of fast-evolving binaries would strongly support the early merger model, offering significant insights into galaxy and SMBH redshift evolution.
@@ 1 Introduction
Last year, the NANOGrav (Agazie et al., 2023a) and other PTA Collaborations reported evidence for the presence of a nanohertz GWB (Reardon et al., 2023; EPTA Collaboration et al., 2023). The stochastic GWB is primarily expected to be the incoherent superposition of gravitational waves from SMBHBs across the entire universe, although other possibilities exist (Afzal et al., 2023).
Current knowledge suggests that the properties of SMBHs are strongly correlated with those of their host galaxies (McConnell & Ma, 2013; Saglia et al., 2016), and the merger of SMBHs is likely to follow the merger of their host galaxies (Enoki et al., 2004; Sesana, 2013). However, calculations of the GWB based on our local galactic astrophysical observations predict a value that is several times lower than the detected one (Agazie et al., 2023b). This discrepancy is not major, given the uncertainties in galactic observations, one can modestly adjust the modeling parameter values to bridge the gap between predictions and observations. Agazie et al. (2023b) has shown that the observed GWB amplitude could be achieved with short binary hardening timescales, higher galaxy number densities, or higher normalization of the $M_{\mathrm{BH}}-M_{\text{bulge}}$ relation, resulting in more SMBH mergers, particularly involving more massive ones.
The unprecedented capabilities of JWST have revolutionized our understanding of the high-redshift universe. So far, JWST has discovered a considerably large number of high-redshift massive galaxies, challenging our current understanding of galaxy formation and evolution (Barrufet et al., 2024; Urbano Stawinski et al., 2024; Wright et al., 2024). The observed number of galaxies at the high-mass end of the galaxy stellar-mass function suggests a considerable number of massive galaxies already exist at $z\sim 5$ (Gottumukkala et al., 2024; Weibel et al., 2024). JWST observations also indicate that SMBHs are much more massive than what the local $M_{\mathrm{BH}}-M_{\text{bulge}}$ relation expects at high redshift (Yue et al., 2024; Pacucci et al., 2023; Mezcua et al., 2024; Juodžbalis et al., 2024).
@@ 2 Population Models
Our objective is to investigate the characteristics of the SMBHB population capable of generating a GWB consistent with the NANOGrav 15-year dataset. We initiate our analysis by synthesizing the population of SMBHs that already merged in the cosmic history of the universe.
The current detected GW background signal is around a factor of four stronger than the reference values suggested by local galactic observations (Agazie et al., 2023b). In this work, we explore two distinct scenarios to mitigate the gap between the prediction and observations.
@@
We fix the lifetime of binaries to be $\tau_{\rm tot}=0.1$ Gyr for all the mergers. With this simplification, we don’t need to consider the evolution details of the binaries. See the Section 5.2 for more discussion on the effects of this simplification.
2.1 Number Density of Galaxy Merger
@@
This distribution is calculated in terms of the stellar mass of the primary galaxy $m_{\star 1}$, the stellar mass ratio $\left(q_{\star}=m_{\star 2}/m_{\star 1}\leqslant 1\right)$, and the redshift $z$. Because the galaxy and the subsequent SMBH merger span a finite timescale $\left(T_{\text{gal-gal}}+\tau_{\rm tot}\right)$, we distinguish between the initial redshift at which a galaxy pair forms $\left(z^{\prime}=z^{\prime}[t]\right.$ at some initial time $\left.t\right)$ and the redshift at which the system becomes a postmerger remant $\left(z=z\left[t+T_{\text{gal}-\text{gal}}+\tau_{\rm tot}\right]\right)$.
@@
There are two different choices of the parameters for $\psi_{0}$ and $m_{\psi,0}$, one derived from local galaxy observation mean values (Tomczak et al., 2014) and the other used parameters from the best-fit GW-driven decay model in Agazie et al. (2023b). These sets of parameters correspond to the Galaxy Observation Mean Value (GOMV) model and the late merger model, respectively, and they are all described by only a single-Schechter function for each.
@@
The early merger model assumes a more complicated GSMF. Figure 2 shows the distribution characterized by the mean values derived from observations Tomczak et al. (2014) and the distribution with an extra component as we proposed. The GSMFs of the early merger model are the sum of the original and added distributions. Equivalently, the early merger model describes the GSMF by a double-Schechter function, with an additional single-Schechter function describing the possible high-redshift massive component of the whole galaxy populations,
@@
Here $\Psi_{0,a}$, $M_{\psi,a}$, $\alpha_{\psi,a}$, $\Psi_{0,o}$, $\alpha_{\psi,o}$, and $M_{\psi,o}$ are also defined in a similar way as Equation 4. $\Psi_{0,o}$, $\alpha_{\psi,o}$, and $M_{\psi,o}$ denote the original component, with values following the same values as GOMV. $\Psi_{0,a}$, $\alpha_{\psi,a}$, and $M_{\psi,a}$ are parameters describing the newly added component in Fig 2 with numerical values listed in Table 1. The values are chosen to produce the observed GWB amplitude.
Although we include an additional population of high-redshift massive galaxies, it turns out that the new GSMF is overall consistent with the local galactic observations (Tomczak et al., 2014). As shown in Figure 3, the galaxy mass distribution of the early merger model is within the uncertainty area for redshifts lower than two, and only slightly deviates from the observations at $z=3$. Therefore, there are no significant disagreements between the early merger model and current observational datasets.
Figure 3: GSMF $\Psi(m,z)$ for the early merger model over the redshift range $30<z<$ 3, following the conventions established in Figure 2. The blue and red shaded areas reflect the parameter values and uncertainties derived from observations, as reported by Tomczak et al. (2014). The early merger model’s GSMF aligns closely with observed data up to $z=3$, the observational limit, demonstrating only minor deviations at this boundary and maintaining full consistency for $z<2$. This alignment indicates no significant discrepancies between the early merger model and current observations.
@@
In practice, we parameterize $P\left(m_{\star 1},q_{\star},z^{\prime}\right)$ and $T_{\text{gal-gal }}\left(m_{\star 1},q_{\star},z^{\prime}\right)$ as redshift-dependent power laws of $m_{\star 1}$, $q_{\star}$, and $z$ following (Chen et al., 2019):
@@ 2.2 SMBH Merger Density
Note that only a fraction of the galaxy stellar mass is in the stellar bulge component $\left(M_{\text{bulge }}=f_{\star,\text{bulge}}\cdot m_{\star}\right)$, which we take to be $f_{\star,\text{bulge}}=0.615$ based on empirical bulge fraction measurements of massive galaxies from Lang et al. (2014) and Bluck et al. (2014). Using the $M_{\mathrm{BH}}-M_{\text{bulge}}$ relationship, we transform the number density of galaxy mergers to a number density of SMBH binaries via
@@ 2.3 SMBHB Population
Notice that the interactions of the binary with the environment are important for the decay of the orbits before the GW emission becomes dominant. These interactions are necessary for the binary to merge within the Hubble time. However, taking into account the details of the interaction is too complicated and beyond the scope of this paper. Depending on the mechanism of interactions, the hardening timescale varies dramatically, so does the population of the binaries. However, we expect the first detected single binary by PTAs to have already entered the GW-dominated phase and is, therefore, insensitive to these interactions. More discussion on this point in Section 5.2.
@@
We generate multiple realizations by drawing many times from a Poisson distribution centered at that value. The frequency bins are typically chosen based on the preference of the data for various red-noise processes, with bins centered at $f_{i}=i/T_{obs}$, where $i=1,\ldots 14$ and $T_{obs}=16.03$ yr (Agazie et al., 2023a).
@@ 3.1 From Populations to GWB
with $q\equiv m_{2}/m_{1}(<1)$ representing the binary mass ratio. Here we assume the mass ratio between two SMBHs is not extreme for the validity of Newtonian approximation. Combining the above equations we have
@@ 3.2 Late Merger vs Early Merger Model
The results are shown in Fig 4, where we compare the characteristic gravitational wave strains, $h_{c}$, as predicted by GOMV, late merger, and early merger models. Grey violin plots illustrate the posterior distributions of $h_{c}$ across frequency bins, derived from an analysis that includes HD-correlated noise and fits for multiple noise spectra (HD-w/MP+DP+CURN), with only the first nine posteriors displayed for clarity.
It is obvious from Fig 4 that the characteristic strain generated by the GOMV (brown dash line) is significantly lower than the observed GWB. The red squares and blue crosses represent the best-fitted power-law spectra for the late and the early merger models based on the median amplitudes of the posteriors from the HD-w/MP+DP+CURN model in Agazie et al. (2023a).
@@ 4.1 Spectral Noise Density
To evaluate the probability of detecting single sources, it is necessary first to understand the noise characteristics of the detector. We utilize the effective strain noise power spectral density, denoted as $S_{\mathrm{eff}}$, which incorporates all aspects of detector noise characterization into a single figure of merit.
@@ 4.2 Evolving Binary & SNR
4.3 Detection Probability of Single Binary
@@ 4.4 Properties of the First Detected Source
Figure 10: PDF of the mass of the loudest binaries for both late and early merger models. The logarithmic scale has been properly accounted for to ensure the area under the PDF represents the correct fraction. Only binaries with SNR $>$ 3 for $T_{\mathrm{span}}=25$ years are presented. Across all models and configurations, a pronounced peak is observed at $M\approx 5\times 10^{9}M_{\odot}$. Notably, the early merger model exhibits an additional peak at higher total masses, originating from the high-redshift massive merger component.
Figure 10 shows the PDF of the mass of the loudest binaries. Across all models and configurations, a prominent peak appears at $M\approx 5\times 10^{9}M_{\odot}$, consistent with the masses of binaries that strongly dominate the GWB signal as reported by Agazie et al. (2023b). Similar to Figure 7, the early merger model exhibits an additional peak at higher total masses. The additional peak originates from the high-redshift massive merger component, and the fraction of very massive binaries ($M\gtrsim 10^{10}M_{\odot}$) aligns with the fraction of high-redshift binaries in Figure 9.
Additionally, the status quo configuration exhibits a slight shift towards higher masses and a significant reduction in the overall count of loud binaries. These features result from the elevated noise spectral density in the status quo configuration, which permits the detection of only very massive binaries from the high-redshift component. Nonetheless, the overall distribution features remain largely insensitive to the assumed PTA.

### v2

# High-Redshift Merger Model for Low-Frequency Gravitational Wave Background

@@ Abstract
In 2023, the Pulsar Timing Array (PTA) Collaborations announced the discovery of a gravitational wave background (GWB), predominantly attributed to supermassive black hole binary (SMBHB) mergers. However, the detected GWB is several times stronger than the default value expected from galactic observations at low and moderate redshifts. Recent findings by the James Webb Space Telescope (JWST) have unveiled a substantial number of massive, high-redshift galaxies, suggesting more massive SMBHB mergers at these early epochs. Motivated by these findings, we propose an “early merger” model that complements the standard merger statistics by incorporating these early, massive galaxies. We compare the early and standard “late merger” models, which assume peak merger rates in the local Universe, and match both merger models to the currently detected GWB. Our analysis shows that the early merger model has a significantly lower detection probability for single binaries and predicts a $\sim 30\%$ likelihood that the first detectable single source will be highly redshifted and remarkably massive with rapid frequency evolution. In contrast, the late merger model predicts a nearly monochromatic first source at low redshift. The future confirmation of an enhanced population of massive high-redshift galaxies and the detection of fast-evolving binaries would strongly support the early merger model, offering significant insights into the evolution of galaxies and SMBHs.
@@ 1 Introduction
Last year, the North American Nanohertz Observatory for Gravitational waves (NANOGrav), the European PTA (EPTA), the Indian PTA (InPTA), the Australia-based Parkes PTA (PPTA), and the Chinese PTA reported evidence for the presence of a nanohertz GWB (Agazie et al., 2023a; EPTA Collaboration et al., 2023; Reardon et al., 2023; Xu et al., 2023). The stochastic GWB is primarily expected to be the incoherent superposition of gravitational waves from SMBHBs across the entire Universe, although other possibilities exist (Afzal et al., 2023).
Current knowledge suggests that the properties of SMBHs are strongly correlated with those of their host galaxies (McConnell & Ma, 2013; Saglia et al., 2016), and the merger of SMBHs is likely to follow the merger of their host galaxies (Enoki et al., 2004; Sesana, 2013a). However, calculations of the GWB based on our local galactic astrophysical observations predict a value that is several times lower than the detected one (Chen et al., 2020; Siwek et al., 2020; Izquierdo-Villalba et al., 2022; Agazie et al., 2023b; Matt et al., 2023; Curyło & Bulik, 2024; EPTA Collaboration et al., 2024). This discrepancy is not major, given the uncertainties in galactic observations, one can modestly adjust the modeling parameter values to bridge the gap between predictions and observations. Agazie et al. (2023b) have shown that the observed GWB amplitude could be achieved with short binary hardening timescales, higher galaxy number densities, or higher normalization of the $M_{\mathrm{BH}}-M_{\text{bulge}}$ relation, resulting in more SMBH mergers, particularly involving more massive ones.
The unprecedented capabilities of JWST have revolutionized our understanding of the high-redshift Universe. So far, JWST has discovered a considerably large number of high-redshift massive galaxies, challenging our current understanding of galaxy formation and evolution (Barrufet et al., 2024; Urbano Stawinski et al., 2024; Wright et al., 2024). The observed number of galaxies at the high-mass end of the galaxy stellar-mass function suggests a considerable number of massive galaxies already exist at $z\sim 5$ (Gottumukkala et al., 2024; Weibel et al., 2024). JWST observations also indicate that SMBHs are much more massive than what the local $M_{\mathrm{BH}}-M_{\text{bulge}}$ relation expects at high redshift (Maiolino et al., 2023; Harikane et al., 2023; Yue et al., 2024; Pacucci et al., 2023; Mezcua et al., 2024; Juodžbalis et al., 2024).
@@ 2 Population Models
Our objective is to investigate the characteristics of the SMBHB population capable of generating a GWB consistent with the NANOGrav, EPTA+InPTA, and PPTA recent results. We initiate our analysis by synthesizing the population of SMBHs that already merged in the cosmic history of the Universe.
The currently detected GWB signal is several times stronger than the reference values suggested by local galactic observations (Chen et al., 2020; Siwek et al., 2020; Izquierdo-Villalba et al., 2022; Agazie et al., 2023b; Matt et al., 2023; Curyło & Bulik, 2024; EPTA Collaboration et al., 2024). In this work, we explore two distinct scenarios to mitigate the gap between the prediction and observations.
@@
We fix the lifetime of binaries to be $\tau_{\rm tot}=0.1$ Gyr for all the mergers. With this simplification, we do not need to consider the evolution details of the binaries. See the Section 5.2 for more discussion on the effects of this simplification.
2.1 Number Density of Galaxy Mergers
@@
This distribution is calculated in terms of the stellar mass of the primary galaxy $m_{\star 1}$, the stellar mass ratio $\left(q_{\star}=m_{\star 2}/m_{\star 1}\leqslant 1\right)$, and the redshift $z$. Because the galaxy and the subsequent SMBH merger span a finite timescale $\left(T_{\text{gal-gal}}+\tau_{\rm tot}\right)$, we distinguish between the initial redshift at which a galaxy pair forms $\left(z^{\prime}=z^{\prime}[t]\right.$ at some initial time $\left.t\right)$ and the redshift at which the system becomes a postmerger remnant $\left(z=z\left[t+T_{\text{gal}-\text{gal}}+\tau_{\rm tot}\right]\right)$.
@@
There are two different choices of the parameters for $\psi_{0}$ and $m_{\psi,0}$, one derived from local galaxy observation mean values (Tomczak et al., 2014) and the other used results of the Phenom+Astro model from Agazie et al. (2023b). These sets of parameters correspond to the Galaxy Observation Mean Value (GOMV) model and the late merger model, respectively, and they are all described by only a single-Schechter function for each.
@@
The early merger model assumes a more complicated GSMF. Figure 2 shows the distributions characterized by the GOMV (Original) and the distribution of an extra component as we proposed (Added). The GSMFs of the early merger model are the sum of the original and added distributions. Equivalently, the early merger model describes the GSMF by a double-Schechter function, with an additional single-Schechter function describing the possible high-redshift massive component of the whole galaxy populations,
@@
Here $\Psi_{0,a}$, $M_{\psi,a}$, $\alpha_{\psi,a}$, $\Psi_{0,o}$, $\alpha_{\psi,o}$, and $M_{\psi,o}$ are also defined in a similar way as Equation 4. $\Psi_{0,o}$, $\alpha_{\psi,o}$, and $M_{\psi,o}$ denote the original component, with values following the same values as GOMV. $\Psi_{0,a}$, $\alpha_{\psi,a}$, and $M_{\psi,a}$ are parameters describing the newly added component in Figure 2 with numerical values listed in Table 1. The values are chosen to produce the observed GWB amplitude.
Although we include an additional population of high-redshift massive galaxies, the new GSMF remains consistent with local galactic observations (Tomczak et al., 2014) for $z\lesssim 2$. In Figure 3, the shaded areas indicate the uncertainties in the GSMFs. These uncertainties are based on the posterior width of the turnover mass $M_{\psi}$, as inferred from Tomczak et al. (2014) using the model described by Equations (3) and (4). We focus on the turnover mass since it dominates the critical high-mass end of the GSMFs.
The GSMF of the early merger model shows a prominent excess at $m\sim 10^{11}M_{\odot}$ for $z>2$. In contrast, the GSMF of the late merger model has moderate deviations from the local galactic observations across all masses and redshifts. These discrepancies can be mitigated by adjusting either the scaling relation or the redshift dependence of the parameters to produce a similar GWB level. The early merger model is more flexible, as its two components can be tuned separately.
However, if we assume a GSMF and scaling relation within local observational error margins and extrapolate to higher redshifts, the resulting GWB is significantly lower than detected. The primary aim of this paper is to investigate the different predictions arising from two distinct models for the SMBHB merger history. Therefore, we do not focus on the specific parameter values within these models and leave detailed modeling for future studies as more observational data become available.
Figure 3: GSMF $\Psi(m,z)$ for various models across the redshift range $31<z<$ 3. Dashed lines represent the GSMF according to the GOMV model, solid lines show the GSMF of the early merger model, and dash-dot lines indicate the late merger model. The shaded areas denote uncertainties due to turnover mass variations, with the uncertainty range inferred from Tomczak et al. (2014). The early merger model’s GSMF closely matches observations up to $z=2$ but diverges notably at $z=3$, exhibiting a prominent excess at $m\sim 10^{11}M_{\odot}$. In contrast, the GSMF of the late merger model shows moderate deviations across all masses and redshifts.
@@
In practice, we parameterize $P\left(m_{\star 1},q_{\star},z^{\prime}\right)$ and $T_{\text{gal-gal }}\left(m_{\star 1},q_{\star},z^{\prime}\right)$ as power laws of $q_{\star}$ and $z$ following Chen et al. (2019):
@@ 2.2 SMBH Merger Density
Note that only a fraction of the galaxy stellar mass is in the stellar bulge component $\left(M_{\text{bulge}}=f_{\star,\text{bulge}}\cdot m_{\star}\right)$, which we take to be $f_{\star,\text{bulge}}=0.615$ based on empirical bulge fraction measurements of massive galaxies from Lang et al. (2014) and Bluck et al. (2014). In principle, the $f_{\star,\text{bulge}}$ could depend on the stellar mass (Moffett et al., 2016), but it is also likely related to the redshift and galaxy type (Huertas-Company et al., 2024; Jin et al., 2024). However, in our study, $f_{\star,\text{bulge}}$ is degenerate with $\mu$ and $\alpha_{\mu}$, and none of these parameters are yet well understood or constrained at high redshifts. Therefore, to simplify the model, we fixed $f_{\star,\text{bulge}}$ as a constant throughout the paper, leaving more detailed treatments for future work.
Using the $M_{\mathrm{BH}}-M_{\text{bulge}}$ relationship, we transform the number density of galaxy mergers to a number density of SMBH binaries via
@@ 2.3 SMBHB Population
Notice that the interactions of the binary with the environment are important for the decay of the orbits before the GW emission becomes dominant. These interactions are necessary for the binary to merge within the Hubble time (EPTA Collaboration et al., 2024; Agazie et al., 2023b). However, taking into account the details of the interaction is too complicated and beyond the scope of this paper. Depending on the mechanism of interactions, the hardening timescale varies dramatically, so does the population of the binaries. However, we expect the first detected single binary by PTAs to have already entered the GW-dominated phase and is, therefore, insensitive to these interactions. More discussion on this point in Section 5.2.
@@
We generate multiple realizations by drawing many times from a Poisson distribution centered at that value. The frequency bins are chosen to align with the data preferences for various red-noise processes in the NANOGrav 15-year dataset (Agazie et al., 2023c). These bins are centered at $f_{i}=i/T_{obs}$, where $i=1,\ldots 14$ and $T_{obs}=16.03$ yr (Agazie et al., 2023a).
@@ 3.1 From Populations to GWB
with $q\equiv m_{2}/m_{1}(\leq 1)$ representing the binary mass ratio. Here we assume the mass ratio between two SMBHs is not extreme for the validity of Newtonian approximation. Combining the above equations we have
@@ 3.2 Late Merger vs Early Merger Model
The results are shown in Fig 4, where we compare the characteristic gravitational wave strains, $h_{c}$, as predicted by GOMV, late merger, and early merger models. Grey violin plots illustrate the posterior distributions of $h_{c}$ across frequency bins, derived from an analysis that includes HD-correlated noise and fits for multiple noise spectra (HD-w/MP+DP+CURN; Agazie et al. 2023a), with only the first nine posteriors displayed for clarity.
Henceforth, we focus on the NANOGrav results, given that EPTA+InPTA and PPTA have detected similar GWB amplitudes, and our subsequent analysis uses pulsar data from NANOGrav. It is obvious from Fig 4 that the characteristic strain generated by the GOMV (brown dash line) is significantly lower than the observed GWB. The red squares and blue crosses represent the best-fitted power-law spectra for the late and the early merger models based on the median amplitudes of the posteriors from the HD-w/MP+DP+CURN model in Agazie et al. (2023a). Compared to the late merger model, the GOMV model features smaller normalizations of the GSMF and the scaling relation, as well as a lower turnover mass for the GSMF. These characteristics result in SMBHB mergers with lower mass and fewer occurrences in the GOMV model. Consequently, the $h_{c}$ produced by the GOMV model is significantly smaller than that of the late merger model. The early merger model, however, includes an additional population of galaxies and therefore an additional population of SMBHB mergers, leading to a higher GWB compared to the GOMV model.
@@ 4.1 Spectral Noise Density
To evaluate the probability of detecting single sources, it is necessary first to understand the noise characteristics of the detector. This study focuses on pulsar data from the NANOGrav 15-year dataset, utilizing the effective strain noise power spectral density ($S_{\mathrm{eff}}$), which incorporates all aspects of detector noise characterization into a single figure of merit.
@@ 4.2 Evolving Binary & SNR
4.3 Detection Probability of a Single Binary
@@ 4.4 Properties of the First Detected Source
Figure 10: PDF of the total mass of the loudest binaries for both late and early merger models. The logarithmic scale has been properly accounted for to ensure the area under the PDF represents the correct fraction. Only binaries with SNR $>$ 3 for $T_{\mathrm{span}}=25$ years are presented. Across all models and configurations, a pronounced peak is observed at $M\approx 5\times 10^{9}M_{\odot}$. Notably, the early merger model exhibits an additional peak at higher total masses, originating from the high-redshift massive merger component.
Figure 10 shows the PDF of the total mass of the loudest binaries. Across all models and configurations, a prominent peak appears at $M\approx 5\times 10^{9}M_{\odot}$, consistent with the masses of binaries that strongly dominate the GWB signal as reported by Agazie et al. (2023b). Similar to Figure 7, the early merger model exhibits an additional peak at higher total masses. The additional peak originates from the high-redshift massive merger component, and the fraction of very massive binaries ($M\gtrsim 10^{10}M_{\odot}$) aligns with the fraction of high-redshift binaries in Figure 9.
Additionally, the status quo configuration exhibits a slight shift towards higher total masses and a significant reduction in the overall count of loud binaries. These features result from the elevated noise spectral density in the status quo configuration, which permits the detection of only very massive binaries from the high-redshift component. Nonetheless, the overall distribution features remain largely insensitive to the assumed PTA.